Find the vectors that join the center of a clock to the hours 1:00, 2:00, and 3:00. Assume the clock is circular with a radius of 1 unit


1
Expert's answer
2021-08-10T06:52:10-0400

Consider the clock given in the figure shows.




The angle made by vectors joining the centre and 1, 2, and 3 and the positive x-axis are as follows:

OA with x-axis=60o

OB with x-axis=30o

OC with x-axis=00

vector OA ="1cos60^oi+1sin60^oj"

at 1 we have "1/2i+\sqrt{3}/2j"

vector OB ="1cos30^oi+1sin60^oj"

at 2 we have "\sqrt{3}/2i" "+1\/2j"

vector OC ="1cos0 i+1sin0 j=i+0j=i" (at 3:00)